OpenStudy (anonymous):

i need this answerd: -(3/2)x=-(2/5)y-(10/3) Rewrite the equation below into standard form.

OpenStudy (texaschic101):

-3/2x = -2/5y - 10/3 (multiply the entire equation by the LCD, which is 30) -45x = - 12y - 100 (add 12y to both sides) -45x + 12y = -100 (multiply by -1 to make x positive) 45x - 12y = 100 <== standard form

OpenStudy (anonymous):

thank you what class are you taking? message me!

OpenStudy (texaschic101):

I do not take classes at this time. I am currently out of school.

OpenStudy (texaschic101):

Is that not an answer choice ?

OpenStudy (anonymous):

no they all equal 12 or 4

OpenStudy (anonymous):

for example 3x-6y=12 thats how they are

OpenStudy (texaschic101):

did I write the problem correctly ?

OpenStudy (anonymous):

3x-6y=12, 3x-6y=-12, -x+2y=4, x-2y=-4 these are my choices!

OpenStudy (texaschic101):

-3/2x = -2/5y - 10/3.......Is this the question....make sure

OpenStudy (texaschic101):

hello ??

OpenStudy (anonymous):

sorry i have a new problem took to long to answer lol.

OpenStudy (texaschic101):

lets see it

OpenStudy (anonymous):

heres the new question: A small child is at a candy store, eager to spend his $1 allowance on candy. His two favorite candies are priced at $0.05 and $0.10, respectively. If the child is intent on spending his entire allowance on candy, which of the following represents a standard form equation that describes this situation?

OpenStudy (texaschic101):

.05x + .10y = 1.00

OpenStudy (texaschic101):

or 5x + 10y = 100

OpenStudy (anonymous):

Match a standard form equation with the graph below.